40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Apr 26 19:35:16 2019, Jon, Update, SUS, New c1susaux installed, passed first round of scripted testing IMG_3254.jpgIMG_3256.jpg
    Reply  Fri May 3 12:48:40 2019, gautam, Update, SUS, 1X4/1X5 cable admin D38CC485-1EB6-4B34-9EB1-2CB1E809A21A.jpeg
       Reply  Tue Jul 2 12:30:53 2019, gautam, Update, Electronics, Acromag crate switched to Sorensens 1X5Sorensens.pdfCDS_20190702.png
          Reply  Tue Jul 2 16:57:09 2019, gautam, Update, CDS, c1sus is flaky c1sus_timing.png
Message ID: 14581     Entry time: Fri Apr 26 19:35:16 2019     Reply to this: 14592
Author: Jon 
Type: Update 
Category: SUS 
Subject: New c1susaux installed, passed first round of scripted testing 

[Jon, Gautam]

Today we installed the c1susaux Acromag chassis and controller computer in the 1X4 rack. As noted in 14580 the prototype Acromag chassis had to first be removed to make room in the rack. The signal feedthroughs were connected to the eurocrates by 10' DB-37 cables via adapters to 96-pin DIN.

Once installed, we ran a scripted set of suspension actuation tests using PyIFOTest. BS, PRM, SRM, MC1, MC2, and MC3 all passed these tests. We were unable to test ITMX and ITMY because both appear to be stuck. Gautam will shake them loose on Monday.

Although the new c1susaux is now mounted in the rack, there is more that needs to be done to make the installation permanent:

  • New 15V and 24V power cables with standard LIGO connectors need to be run from the Sorensenn supplies in 1X5. The chassis is currently powered by bench supplies sitting on a cart behind the rack.
  • All 24 new DB-37 signal cables need to be labeled.
  • New 96-pin DIN connectors need to be put on two ribbon cables (1Y5_80 B, 1Y5_81) in the 1X4 rack. We had to break these connectors to remove them from the back of the eurcrates.
  • General cleanup of any cables, etc. left around the rack. We cleaned up most things this evening.
  • Rename the host computer c1susaux2 --> c1susaux, and update the DNS lookup tables on chiara.

On Monday we plan to continue with additional scripted tests of the suspensions.


gautam - some more notes:

  • Backplane connectors for the SUS PD whitening boards, which now only serve the purpose of carrying the fast BIO signals used for switching the whitening, were moved from the P1 connector to P2 connector for MC1, MC2, MC3, ITMX, ITMY, BS and PRM.
  • In the process, the connectors for BS and PRM were detatched from the ribbon cable (there wasn't any good way to unseat the connector from the shell that I know of). These will have to be repaired by Chub, and the signal integrity will have to be checked (as they have to be for the connectors that are allegedly intact).
  • While we were doing the wiring, I disconnected the outputs of the coil driver board going to the satellite box (front panel DB15 connector on D010001). These were restored after our work for the testing phase.
  • The backplane cables to the eurocrate housing the coil driver boards were also disconnected. They are currently just dangling, but we will have to clean it up if the new crate is performing alright.
  • In general the cable routing cleanliness has to be checked and approved by Chub or someone else qualified. In particular, the power leads to the eurocrate are in the way of the DIN96-DB37 adaptor board of Johannes' design, particularly on the SUS PD eurocrate.
  • Tapping new power rails for the Acromag chassis will have to be done carefully. Ideally we shouldn't have to turn off the Sorensens.
  • There are some software issues we encountered today connected with the networking that have to be understood and addressed in a permanent way.
  • Sooner rather than later, we want to reconnect the Acromag crate that was monitoring the PSL channels, particularly given the NPRO's recent flakiness.
  • The NPRO was turned back on (following the same procedure of slowly dialing up the injection current). Primary motivation to see if the mode cleaner cavity could be locked with the new SUS electronics. Looks like it could. I'm leaving it on over the weekend...
Attachment 1: IMG_3254.jpg  1.976 MB  Uploaded Fri Apr 26 20:36:25 2019  | Hide | Hide all
IMG_3254.jpg
Attachment 2: IMG_3256.jpg  1.651 MB  Uploaded Fri Apr 26 20:36:45 2019  | Hide | Hide all
IMG_3256.jpg
ELOG V3.1.3-